A low Earth orbit (LEO) is an orbit around Earth with an altitude between 160 kilometers (99 mi) (orbital period of about 88 minutes). If the radius of the earth is around 6400km, what is the speed of a satellite (in Km/s) on this LEO?
Answer #1

Radius of Earth = RE = 6400 km

Altitude of the satellite = H = 160 km

Radius of orbit of the satellite = R

R = RE + H

R = 6400 + 160

R = 6560 km

Orbital period of the satellite = T = 88 min = 88 x (60) sec = 5280 sec

Orbital speed of the satellite = V

VT = 2R

V(5280) = 2(6560)

V = 7.806 km/s

Speed of the satellite = 7.806 km/s